OpenStudy (anonymous):

PLEASE HELP dy/dx+ 2xy + xy^4 = 0

OpenStudy (amistre64):

burnoodle

OpenStudy (amistre64):

\[y' + 2xy = -xy^4\] \[(y' + 2xy = -xy^4)y^{-4}\] \[y^{-4} y' + 2xy^{-3 } = -x\]

OpenStudy (anonymous):

Not sure im getting it

OpenStudy (anonymous):

y^-4? whres that coming from

OpenStudy (amistre64):

substitute a new funtion of y in place of y; \[\begin{array}\\ z=y^{-3}\\\\ z^{-1/3}=y& :z^{4/3}=y^{-4}\\\\ -\frac{1}{3}z^{-4/3}z' =y' \end{array} \] \[y^{-4} y' + 2xy^{-3 } = -x\] \[z^{4/3} (-\frac{1}{3}z^{-4/3}z') + 2x\ z = -x\] \[-\frac{1}{3}z' + 2x\ z = -x\]

OpenStudy (amistre64):

it comes from dividing off the y^4 to get all the ys to one side

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

multiply thru by -3 to clear our z' \[-3(-\frac{1}{3}z' + 2x\ z = -x)\] \[z' + (-6x)\ z = 3x\] and we are to a form that we should be familiar with now

OpenStudy (amistre64):

solve for z and then revert back to y z = (.......) = y^-3

OpenStudy (anonymous):

wow..you good..i need to look at it a few mins to understand

OpenStudy (amistre64):

thanx, if you have questions, feel free to ask

OpenStudy (anonymous):

how do i find an intersection of 2 vectors

OpenStudy (anonymous):

the question actualy says find the S intesect Tand hence find a basis and the dimention, where S={(a+b,a,-a,-b)} and T={(x,2y,-y,-x)}

OpenStudy (amistre64):

IF they intersect, then there component parts will be in the same place at the same time

OpenStudy (amistre64):

S=: 1a+1b | 1| | 1| 1a+0b = a | 1|+ b | 0| -1a+0b |-1| | 0| 0a-1b | 0| |-1| all linear combinations of the span of independant vectors form a vector space. \[col\ S=span\left\{ \begin{pmatrix} 1\\1\\-1\\0 \end{pmatrix}, \begin{pmatrix}1\\0\\0\\-1 \end{pmatrix}\right\}\] T=: 1x+0y | 1| | 0| 0x+2y = x | 0|+ y | 2| 0x-1y | 0| |-1| -1x+0y |-1| | 0| same reasoning as for S \[col\ T=span\left\{ \begin{pmatrix}1\\0\\0\\-1 \end{pmatrix}, \begin{pmatrix}0\\2\\-1\\0 \end{pmatrix}\right\}\]

OpenStudy (amistre64):

I notice that the span of each vector space contains the same vector; [1,0,0,-1] so this is at least a line of intersection; for when a=0 and y=0 and b=x; we have then same object defined

OpenStudy (amistre64):

2 vectors will span at most a plane; the planes themselves stretch thru R^4, so these are a subspace of R^4 ..

OpenStudy (amistre64):

the other 2 vectors, if they form equal ratios from their component parts; they are parallel; for any vector that is a scalar multiple of another retains the same ratio of parts 2 3 5 8 14 21 35 56 ----------------- 1/7 1/7 1/7 1/7 0,2,-1,0 1,1,-1,0 -------- 0,2, 1,und ; since their ratios are not consistent they are not scalar multiples and therefore NOT parallel